OpenStudy (anonymous):

how to find the value of x in a right triangle. I have the 2 legs..need to find the hyptonuse? is it just the pythagourm theroum?

OpenStudy (saifoo.khan):

\[b^2 + p^2 = h^2\]b=base p=perpendicular, h=hypotensue,

OpenStudy (saifoo.khan):



OpenStudy (anonymous):

yes, you just use pythagoras' theorem :)

OpenStudy (saifoo.khan):

Hey, salam zara! :D

OpenStudy (anonymous):

thanks! sorry about the spelling!

OpenStudy (anonymous):

Haha, don't worry :)